2dy/dx-2y=x5sin2x-x3+4x4

1
Expert's answer
2021-01-19T01:48:23-0500

To solve linear equation :

"2\frac{dy}{dx}-2y=x^5\sin{2x}-x^3+4x^4"

Lets divide both sides of equation by 2 and multiply both sides by "f(x)=e^{\int-1dx'}=e^{-x}"

"e^{-x}\frac{dy}{dx}-e^{-x}y=\frac{e^{-x}}{2}(x^5\sin{2x}-x^3+4x^4)"

Lets substitute "-e^{-x}=\frac{d}{dx}(e^{-x})"

"e^{-x}\frac{dy(x)}{dx}+\frac{d}{dx}(e^{-x})y(x)=\frac{1}{2}e^{-x}(x^5\sin{2x}-x^3+4x^4)"

Apply the reverse product rule "f\frac{dg}{dx}+g\frac{df}{dx}=\frac{d}{dx}(fg)" to left-hand side:

"\frac{d}{dx}(e^{-x}y(x))= \frac{1}{2}e^{-x}(x^5\sin{2x}-x^3+4x^4)"

Integrate both sides with respect to x:

"\int\frac{d}{dx}(e^{-x}y(x))dx= \int\frac{1}{2}e^{-x}(x^5\sin{2x}-x^3+4x^4)dx"

Evaluate the integrals:

"e^{-x}y(x)= \frac{1}{6250}e^{-x}(-3125(4x^4+15x^3+45x^2+90x+90)-2\cos(2x)(625x^5+1250x^4-500x^3-3600x^2-2280x+528)+(-625x^5+1875x^4+5500x^3+2100x^2-4920x-2808)\sin(2x))+c_1"

by dividing both sides by "e^{-x}"

"y(x)= \frac{e^x}{6250}e^{-x}(-3125(4x^4+15x^3+45x^2+90x+90)-2\cos(2x)(625x^5+1250x^4-500x^3-3600x^2-2280x+528)+(-625x^5+1875x^4+5500x^3+2100x^2-4920x-2808)\sin(2x))+c_1"